# Hopping Too Late: Exploring the Limitations of Large Language Models on Multi-Hop Queries

### v1

@@ 6 Related Work
Tracing and locating factual knowledge has been the focus of many recent works aiming to undercover and control the way transformers make predictions. Several of these works are motivated by knowledge editing (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022; Zhang et al., 2024) while others focus on understanding the inner workings of transformers (Dai et al., 2022; Geva et al., 2022, 2023). Our work follows and expands on this second group by extending them to a more complex case, two-hop queries.
The question whether models can implicitly reason over stored parametric knowledge has also gained recent popularity. One avenue of related research includes circuit discovery (Nanda et al., 2023; Wang et al., 2023b; Conmy et al., 2023; Brinkmann et al., 2024), but these works focus mainly on synthetic tasks. More closely related to this paper, recent works (Sakarvadia et al., 2023; Yang et al., 2024; Li et al., 2024a) have found existential evidence of latent reasoning in multi-hop queries, albeit without accounting for the flow of information throughout layers and positions. In a concurrent work by Wang et al. (2024) a multi-hop reasoning circuit similar to ours was discovered in a small transformer language model trained on synthetic data until grokking occurs. Our work finds further proof of such latent reasoning in a more practical setting that includes large language models, real world knowledge and a regular pretraining setup. We additionally use different mechanistic interpretability methods combined with an understanding of the information flow during the forward pass.
Even without fully identifying the mechanism behind latent reasoning, several works have tried to address models shortcomings. One such work by Sakarvadia et al. (2023) assumes knowledge of the bridge entity, which is then injected directly into the computation. Likewise, recent work by Li et al. (2024a) also assumes knowledge of the bridge entity and uses knowledge editing techniques to improve performance, which is known to have potentially disruptive effects on models (Zhang et al., 2024; Hsueh et al., 2024; Li et al., 2024b). Both these previously proposed methods have drawbacks, which a potential method based on back-patching would not share.
@@ 7 Conclusion
Our work examines only two-hop queries. While we expect our methods and conclusions to hold for queries with three or more hops, and even for unrelated cases of latent reasoning, we leave these areas for future work.
@@ Appendix A Technical Details
We use the Baukit library Bau (2024) for several of the experiments. Each experiment was run on 1-4 A100 or H100 GPUs and lasted at most 24 hours. The models were retrieved and run using the HuggingFace Transformers library Wolf et al. (2020). When generating with sampling we use the default parameters from the Transformers library. We use full precision for the 7B, 8B and 13B models, and half precision for the 70B model.
Appendix B Information Propagates to Last Token
B.1 Method
In similar fashion to previous works (Wang et al., 2023b; Geva et al., 2023), we block hidden representations from attending to other representations at specific layers and test if the model’s final generation changes. We perform the knockout by setting the weight of the attention from a source to a target position at a specific layer to $\infty$- ∞, causing this attention edge to contribute nothing to the residual update performed on the source representation. We experiment with blocking a window of layers as previous work (Geva et al., 2023) shows that information propagation is spread among multiple layers.
B.2 Experiment
We run attention knockout blocking information propagating from $t_{1}$ to $t_{2}$. We block a window of 7 layers in order to capture information flowing through multiple layers. If blocking a set of layers causes the model to no longer correctly predict the answer (in the incorrect setting we check for a change in the generated text), we consider these layers to contain information critical to the prediction.
B.3 Results
Model Setting Detected Mean Layer
LLaMA 2 7B Correct $85.82\%$ $12.83$
Incorrect $95.44\%$ $8.10$
LLaMA 2 13B Correct $81.94\%$ $15.45$
Incorrect $71.11\%$ $16.99$
LLaMA 3 8B Correct $78.36\%$ $10.15$
Incorrect $82.74\%$ $7.45$
LLaMA 3 70B Correct $90.39\%$ $29.14$
Incorrect $94.28\%$ $20.06$
Table 5: Percentages and mean layer of cases where critical information was first detected propagated from $t_{1}$ to $t_{2}$ using attention knockout, projection, or Patchscopes.
Model Setting Resolved Cases Mean Layer
LLaMA 2 7B Correct $41.4\%$ $15.3$
Incorrect $36.7\%$ $16.4$
LLaMA 2 13B Correct $34.2\%$ $18.2$
Incorrect $37.5\%$ $17.9$
LLaMA 3 8B Correct $37.4\%$ $14.3$
Incorrect $25.6\%$ $15.3$
LLaMA 3 70B Correct $46.9\%$ $35.2$
Incorrect $46.2\%$ $34.5$
Table 6: The results of Patchscopes when decoding $e_{2}$ from $t_{2}$. The table reports the percentage of cases where $e_{2}$ was successfully decoded from $t_{2}$. The table additionally reports the mean of the first layer $e_{2}$ was successfully decoded.
Table 6 contains the percentage of cases $e_{2}$ was decoded from $t_{2}$, showing that this position indeed represents the bridge entity in 25%-46% of cases. Comparing between the results of Patchscopes on $t_{2}$ to those on $t_{1}$ (Table 2) we observe a stark difference in the successful source layers. The results show that the $t_{2}$ first encodes $e_{2}$ during the mid-upper layers, while $t_{1}$ does so in the lower layers. This can further be seen in the Patchscopes heat-maps in Figure 7. This gives additional reason to believe that a sequential computation is taking place.
Appendix C Patchscopes First Decoded Layers
@@
(b) LLaMA 3 8B
(c) LLaMA 3 70B
@@
Figure 6 depicts the percentage of cases per layer where target entities were first successfully decoded Patchscopes for additional models (as described in §3).
Appendix D Patchscopes Heat-maps
@@
Figure 7: Patchscopes heat-maps of successful source and target layers for all models.
Figure 7 depicts the heat-maps of Patchscope success cases as introduced in §3. We report Patchscope experiments for decoding $e_{2}$ from $t_{1}$, $e_{2}$ from $t_{2}$, and $e_{3}$ from $t_{2}$ for all models.
Appendix E First Observed Layers of Pathway Stages
@@
(c) LLaMA 3 70B
Figure 8: A comparison of the first layers of each stage in the pathway between correct and incorrect cases for additional models.
Figure 8 depicts a comparison of the first layers of each stage in the pathway between correct and incorrect cases (as shown in §4) for additional models.
Appendix F Back-patching Heat-maps
@@
(c) Source token is $t_{1}$, LLaMA 3 8B
(d) Source token is $t_{2}$, LLaMA 3 8B
(e) Source token is $t_{1}$, LLaMA 3 70B
(f) Source token is $t_{2}$, LLaMA 3 70B
Figure 9: Back-patching heat-maps of successful source and target layers case for additional models.
Figure 9 depicts the back-patching heap-maps of successful source and target layers (as shown in §5) for addition models.

### v2

@@ 7 Related Work
Tracing and locating factual knowledge has been the focus of many recent works aiming to undercover and control the way transformers make predictions. Several of these works are motivated by knowledge editing (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022; Zhang et al., 2024) while others focus on understanding the inner workings of transformers (Dai et al., 2022; Geva et al., 2022, 2023). Our work follows and expands on this second group by extending them to the more complex case of two-hop queries, which require multiple factual knowledge extractions and compositions.
The question whether models can implicitly reason over stored parametric knowledge has also gained recent popularity. One avenue of related research includes circuit discovery (Nanda et al., 2023; Wang et al., 2023b; Conmy et al., 2023; Brinkmann et al., 2024), but these works focus mainly on synthetic tasks. More closely related to this paper, recent works (Sakarvadia et al., 2023; Yang et al., 2024; Li et al., 2024a) have found existential evidence of latent reasoning in multi-hop queries, albeit without accounting for the flow of information throughout layers and positions. In a concurrent work by Wang et al. (2024) a multi-hop reasoning circuit similar to ours was discovered in a small transformer language model trained on synthetic data until grokking occurs. Our work finds further proof of such latent reasoning in a more practical setting that includes large language models, real world knowledge and a regular pretraining setup. We additionally use different mechanistic interpretability methods and account for the flow of information during the forward pass.
Even without fully identifying the mechanism behind latent reasoning, several works have tried to address model’s shortcomings in this area. One such work by Sakarvadia et al. (2023) assumes prior knowledge of the bridge entity, whose hidden representation is then strategically injected directly into the computation. Likewise, recent work by Li et al. (2024a) also assumes knowledge of the bridge entity and then employs knowledge editing techniques to improve performance, which is known to have potentially disruptive effects on models (Zhang et al., 2024; Hsueh et al., 2024; Li et al., 2024b). Both these previously proposed methods have drawbacks, which a potential method based on back-patching would not share.
@@ 8 Conclusion
Our work examines only two-hop queries, although we expect queries with three or more hops and even additional unrelated reasoning tasks to involve similar mechanisms to those we observed. This is due to previous (Wang et al., 2023b; Brinkmann et al., 2024) and concurrent (Wang et al., 2024) works in different (if sometimes small or synthetic) settings observing pathways that can be broken down into concrete computational steps that match intuitive reasoning steps. These results resemble our finding that two-hop questions are broken down and answered sequentially by LLMs. We leave such expansion to additional settings for future work.
@@ Appendix A Technical Details
We use the Baukit library Bau (2024) for several of the experiments. Each experiment was run on 1-4 A100 or H100 GPUs and lasted at most 24 hours. The models were retrieved and run using the HuggingFace Transformers library Wolf et al. (2020). When generating with sampling we use the default parameters from the Transformers library. We use half precision for the 70B model and full precision for all other models.
Appendix B Patchscopes First Decoded Layers
@@
(b) LLaMA 2 13B
(c) LLaMA 3 8B
(d) LLaMA 3 70B
(e) Pythia 6.9B
(f) Pythia 12B
@@
Figure 6 depicts the percentage of cases per layer where target entities were first successfully decoded using Patchscopes for all models (as described in §3).
Appendix C Patchscopes Heat-maps
@@
Figure 7: Patchscopes success heat-maps for LLaMA models.
(a) $e_{2}$ decoded from $t_{1}$, Pythia 6.9B
(b) $e_{2}$ decoded from $t_{2}$, Pythia 6.9B
(c) $e_{3}$ decoded from $t_{2}$, Pythia 6.9B
(d) $e_{2}$ decoded from $t_{1}$, Pythia 12B
(e) $e_{2}$ decoded from $t_{2}$, Pythia 12B
(f) $e_{3}$ decoded from $t_{2}$, Pythia 12B
Figure 8: Patchscopes success heat-maps for Pythia models.
Figures 7 and 8 depict the heat-maps of Patchscope success cases as introduced in §3. We report Patchscope experiments for decoding $e_{2}$ from $t_{1}$, $e_{2}$ from $t_{2}$, and $e_{3}$ from $t_{2}$ for all models.
Appendix D First Observed Layers of Pathway Stages
@@
(c) LLaMA 3 8B
(d) LLaMA 3 70B
(e) Pythia 6.9B
(f) Pythia 12B
Figure 9: A comparison of the first layers of each pathway stage between correct and incorrect cases for all models.
Figure 9 depicts a comparison of the first layers of each stage in the observed pathway between correct and incorrect cases (as shown in §5) for all models.
Appendix E Back-patching Heat-maps
@@
(c) Source token is $t_{1}$, LLaMA 2 13B
(d) Source token is $t_{2}$, LLaMA 2 13B
(e) Source token is $t_{1}$, LLaMA 3 8B
(f) Source token is $t_{2}$, LLaMA 3 8B
(g) Source token is $t_{1}$, LLaMA 3 70B
(h) Source token is $t_{2}$, LLaMA 3 70B
Figure 10: Back-patching heat-maps of successful source and target layers for LLaMA models.
(a) Source token is $t_{1}$, Pythia 6.9B
(b) Source token is $t_{2}$, Pythia 6.9B
(c) Source token is $t_{1}$, Pythia 12B
(d) Source token is $t_{2}$, Pythia 12B
Figure 11: Back-patching heat-maps of successful source and target layers for Pythia models.
Figures 10 and 11 depict the back-patching heap-maps of successful source and target layers (as shown in §6) for all models.